OpenStudy (anonymous):

If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

s= 1/3 A= 18 (1/3)^2 = 18(1/9) =2

OpenStudy (michele_laino):

hint: \[\Large A = 18{s^2} = 18 \times {\left( {\frac{1}{3}} \right)^2} = ...\]

OpenStudy (anonymous):

it is 2